# thanks: Corresponding author

### v1

@@ IV Summary and Concluding Remarks
In this work, we have examined the hybrid variant of the CC($P$;$Q$) methodology introduced in Ref. Gururangan et al., 2021, in which the leading higher–than–doubly excited determinants in the $P$ space are identified, in an automated fashion, using the sequences of Hamiltonian diagonalizations generated with the CIPSI algorithm.Huron et al. (1973); Garniron et al. (2017, 2019) In order to thoroughly test the CIPSI-driven CC($P$;$Q$) formalism and obtain useful insights into its performance, we have focused on recovering the lowest-energy singlet and triplet potentials of cyclobutadiene along its automerization coordinate and the gap between them resulting from the full CCSDT computations which, based on comparisons with the perturbatively corrected and extrapolated CIPSI and DEA-EOMCC(4p-2h)-level data, obtained in this study as well, provide reliable information. To do so, we have constructed an approximate, $D_{\rm 2h}$-symmetric, one-dimensional automerization pathway connecting the rectangular reactant and product species via the square TS structure on the ground-state singlet potential using the information taken from Ref. Eckert-Maksić et al., 2006 and performed a large number of CC($P$;$Q$) calculations for the lowest ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene at the selected nuclear geometries along the resulting path, where for each state and for each geometry, we have explored a wide range of values of the CIPSI wave function termination parameter $N_{\text{det(in)}}$ that controls the Hamiltonian diagonalization sequences preceding the CC($P$) and CC($P$;$Q$) runs. We have demonstrated that the CIPSI-driven CC($P$;$Q$) calculations are capable of accurately approximating the high-level CCSDT energetics of the lowest singlet and triplet states of cyclobutadiene across the entire automerization pathway, to within small fractions of a millihartree for total energies and 0.1–0.2 kcal/mol for the singlet–triplet gaps, using tiny fractions of the triply excited determinants, on the order of 1% of all triples, in the underlying $P$ spaces extracted from the relatively inexpensive CIPSI diagonalizations in spaces that are orders of magnitude smaller than the numbers of cluster amplitudes used by CCSDT. This extraordinary performance of the CIPSI-driven CC($P$;$Q$) approach applies to both the less demanding reactant/product region, where the lowest singlet and triplet states of cyclobutadiene are largely single-configurational and the absolute values of the singlet–triplet gap exceed 30 kcal/mol, and the vicinity of the TS structure on the ground-state singlet potential, where the high-spin triplet state retains its weakly correlated, single-determinantal, nature, but the singlet state, separated from its triplet counterpart by only $\sim$5 kcal/mol, becomes multiconfigurational, strongly correlated, and characterized by large and highly nonperturbative $T_{3}$ correlations, which are strongly coupled to the one- and two-body components of $T$ and which result in failure of CR-CC(2,3) and other noniterative triples corrections to CCSD. Interestingly, the uncorrected CC($P$) computations using similarly compact excitation spaces can be accurate as well, reproducing the CCSDT values of the singlet–triplet gap across the entire automerization pathway to within $\sim$1–2 kcal/mol, improving the poor CCSD and CR-CC(2,3) results in the vicinity of the TS geometry, and reaffirming the usefulness of incorporating the leading triply excited determinants into the underlying $P$ spaces, but the total energies of the lowest ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene obtained with the CC($P$) approach converge to their CCSDT parents with $N_{\text{det(in)}}$ very slowly, so one has to rely on error cancellations to obtain accurate singlet–triplet gaps with CC($P$). The $\delta(\mbox{$P$;$Q$})$ corrections are very helpful in this regard. They reduce errors in the total CC($P$) energies of the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene by orders of magnitude while substantially improving the resulting singlet–triplet gaps, making their convergence toward CCSDT smoother and more systematic. Given the relatively low costs of determining the $\delta(\mbox{$P$;$Q$})$ corrections compared to the preceding CIPSI and CC($P$) steps and the enormous benefits resulting from their application in the CIPSI-driven CC($P$;$Q$) calculations, we recommend using CC($P$;$Q$).

### v2

@@ IV Summary and Concluding Remarks
In this work, we have examined the hybrid variant of the CC($P$;$Q$) methodology introduced in Ref. Gururangan et al., 2021, in which the leading higher–than–doubly excited determinants in the $P$ space are identified, in an automated fashion, using the sequences of Hamiltonian diagonalizations generated with the CIPSI algorithm.Huron et al. (1973); Garniron et al. (2017, 2019) In order to thoroughly test the CIPSI-driven CC($P$;$Q$) formalism and obtain useful insights into its performance, we have focused on recovering the lowest-energy singlet and triplet potentials of cyclobutadiene along its automerization coordinate and the gap between them resulting from the full CCSDT computations which, based on comparisons with the perturbatively corrected and extrapolated CIPSI and DEA-EOMCC(4p-2h)-level data, obtained in this study as well, provide reliable information. To do so, we have constructed an approximate, $D_{\rm 2h}$-symmetric, one-dimensional automerization pathway connecting the rectangular reactant and product species via the square TS structure on the ground-state singlet potential using the information taken from Ref. Eckert-Maksić et al., 2006 and performed a large number of CC($P$;$Q$) calculations for the lowest $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene at the selected nuclear geometries along the resulting path, where for each state and for each geometry, we have explored a wide range of values of the CIPSI wave function termination parameter $N_{\text{det(in)}}$ that controls the Hamiltonian diagonalization sequences preceding the CC($P$) and CC($P$;$Q$) runs. We have demonstrated that the CIPSI-driven CC($P$;$Q$) calculations are capable of accurately approximating the high-level CCSDT energetics of the lowest singlet and triplet states of cyclobutadiene across the entire automerization pathway, to within small fractions of a millihartree for total energies and 0.1–0.3 kcal/mol for the singlet–triplet gaps, using tiny fractions of the triply excited determinants, on the order of 1% of all triples for the cc-pVDZ basis set and 0.1–0.2% for cc-pVTZ, in the underlying $P$ spaces extracted from the relatively inexpensive CIPSI diagonalizations in spaces that are orders of magnitude smaller than the numbers of cluster amplitudes used by CCSDT. This extraordinary performance of the CIPSI-driven CC($P$;$Q$) approach applies to both the less demanding reactant/product region, where the lowest singlet and triplet states of cyclobutadiene are largely single-configurational and the absolute values of the singlet–triplet gap exceed 30 kcal/mol, and the vicinity of the TS structure on the ground-state singlet potential, where the high-spin triplet state retains its weakly correlated, single-determinantal, nature, but the singlet state, separated from its triplet counterpart by only a few kcal/mol, becomes multiconfigurational, strongly correlated, and characterized by large and highly nonperturbative $T_{3}$ correlations, which are strongly coupled to the one- and two-body components of $T$ and which result in failure of CR-CC(2,3) and other noniterative triples corrections to CCSD. Interestingly, the uncorrected CC($P$) computations using similarly compact excitation spaces can be accurate as well, reproducing the CCSDT values of the singlet–triplet gap across the entire automerization pathway to within $\sim$1–2 kcal/mol, improving the poor CCSD and CR-CC(2,3) results in the vicinity of the TS geometry, and reaffirming the usefulness of incorporating the leading triply excited determinants into the underlying $P$ spaces, but the total energies of the lowest $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene obtained with the CC($P$) approach converge to their CCSDT parents with $N_{\text{det(in)}}$ very slowly, so one has to rely on error cancellations to obtain accurate singlet–triplet gaps with CC($P$). The $\delta(\mbox{$P$;$Q$})$ corrections are very helpful in this regard. They reduce errors in the total CC($P$) energies of the $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene by orders of magnitude while substantially improving the resulting singlet–triplet gaps, making their convergence toward CCSDT smoother and more systematic. Given the relatively low costs of determining the $\delta(\mbox{$P$;$Q$})$ corrections compared to the preceding CIPSI and CC($P$) steps and the enormous benefits resulting from their application in the CIPSI-driven CC($P$;$Q$) calculations, we recommend using CC($P$;$Q$).